OpenStudy (anonymous):

A spherical balloon bursts when the surface area(S) exceeds 500 cm^2. If the radius changes at a rate of 23 cm/s, what is the rate of change of volume at the instant the balloon bursts? (V=(4/3)(pi)r^3) (S=4(pi)r^2)

OpenStudy (anonymous):

i would set the surface area equal to 500 cm^2, to find out what the radius is when the balloon will burst: \[500 = 4\pi r^2 \Rightarrow r = \sqrt{\frac{500}{4\pi}}\]

OpenStudy (anonymous):

r=\[\sqrt{125\pi}\] when balloon burst..

OpenStudy (anonymous):

Then, take the derivative of the Volume equation, and plug in that value of r and dr/dt (which is 23 cm/s)

OpenStudy (amistre64):

V=(4/3)(pi)r^3 ; derive the sides dV/dt = (2)(4pi/3) * r^2 * dr/dt ; this gives us the rate of change of the volume with respect to time

OpenStudy (amistre64):

since dr/dt = 23 as stated in the problem; all we have to do is determine the value of "r" to solve right?

OpenStudy (amistre64):

and of course we have to correct for my inherent stupidity :) dV/dt = (3)(4pi/3) * r^2 * dr/dt ; changed the 2 to 3 lol

OpenStudy (anonymous):

So... What you're saying is...?

OpenStudy (amistre64):

it might be interesting to note that the derivative of volume becomes surface area ....

OpenStudy (amistre64):

dV/dt = (3)(4pi/3) * r^2 * dr/dt V' = 4pi r^2 * dr/dt ; and dr/dt = 23 so = 4pi r^2 * 23

OpenStudy (anonymous):

F'(x)=4/3pir^3 ?

OpenStudy (anonymous):

12/9 pi r^2 ?

OpenStudy (amistre64):

since r = sqrt(125/pi) when it bursticates V' = 4pi (sqrt(125/pi))^2 * 23 = 4pi (125/pi)(23) ; the pis cancels = 4(125)(23) ; and whatever that equals in the end

OpenStudy (anonymous):

dv/dt=d(4/3*pi*r^3)/dt=4/3*3*pi*r^2*(dr/dt) =4pi*r^2*(dr/dt)=500*23 The key is the chain rule. I think

OpenStudy (amistre64):

no, there is no chain rule needed in this one

OpenStudy (anonymous):

ohhhh!!! I see. kind of :x Lol. What do you mean by bursticates? ami?(:

OpenStudy (amistre64):

Kr^3 derives to 3Kr^2, no chains required :)

OpenStudy (amistre64):

bursticate, when the bubble pops; like when the housing market took a dive ;)

OpenStudy (anonymous):

ohhhh! okay thankss!

OpenStudy (amistre64):

i think i see what prot was saying; the dr/dt pops out in the end ... yes

OpenStudy (amistre64):

i dont think that is the chain rule perse; Ive always called it a "derived bit"

OpenStudy (amistre64):

its always there, but when its like dx/dx it just goes to 1 and vanishes from teh scene

OpenStudy (anonymous):

Anyway. dv/dt=(dv/dr)*(dr/dt) V=4/3pi*r^3 I think it is the chain rule. But That's ok